OpenStudy (superhelp101):

(a-b+c-d)^2 I know how to foil it, but my teacher wants us to use shortcuts. Such like Perfect square trinomial, binomial cubed, Dots, sum of two cubes, diff of two cubes. I can't seem to find any shortcuts. Any help?

OpenStudy (superhelp101):

@aum

OpenStudy (superhelp101):

@IMStuck @ganeshie8 @dumbcow any help?

ganeshie8 (ganeshie8):

(a-b+c-d)^2 = ( `(a-b)` + `(c-d)` )^2

ganeshie8 (ganeshie8):

use perfect square trinomial formula, expand it out

OpenStudy (superhelp101):

oh okay thanks! :)

ganeshie8 (ganeshie8):

(x+y)^2 = x^2 + 2xy + y^2

OpenStudy (superhelp101):

@aum do you think you can continue from ganeshie ?

OpenStudy (aum):

\[ (a-b+c-d)^2 = (~ (a-b) + (c-d) ~)^2 = (a-b)^2 + 2(a-b)(c-d) + (c-d)^2 = ? \]

OpenStudy (aum):

Think of (a - b) as x and (c - d) as y \[ (a-b+c-d)^2 = (~ (a-b) + (c-d) ~)^2 \\ x = (a-b); ~~~ y = (c-d) \\ (x + y)^2 = x^2 + 2xy + y^2 \\ (a-b)^2 + 2(a-b)(c-d) + (c-d)^2 = ? \]

OpenStudy (superhelp101):

okay that makes sense! Let me solve that and let you know what I got.. just a sec ;)

OpenStudy (superhelp101):

it's a2-2ab+b2+2ac-2ad-2bc+2bd+c2-2cd+d2

OpenStudy (aum):

Correct. You can rearrange the term in a different order or leave it as it is. \(a^2 + b^2 + c^2 + d^2 - 2ab + 2ac - 2ad - 2bc + 2bd + 2cd \)

OpenStudy (superhelp101):

Thank you! :)

OpenStudy (aum):

You are welcome.

OpenStudy (superhelp101):

Oh and also thank you to @ganeshie8 ! You both helped me!